Check my work Wilmington Company has two manufacturing departments--Assembly and Fabrication. It considers all of its manufacTHANK YOU !!
Answer #1

1) Predetermine overhead rate = 11000000/200000 = 55 per labor hour

Manufacturing overhead applied to Job Bravo = 44*55 = $2420

2) Assembly overhead rate = 5250000/125000 = 42 per DLH

Fabrication overhead rate = 5750000/250000 = 23 per MH

Manufacturing overhead applied from Assembly to Job Bravo = 26*42 = 1092

Manufacturing overhead applied from Fabrication to Job Bravo = 21*23 = 483

Total departmental manufacturing overhead applied to Job Bravo = 1092+483 = 1575